OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x^2 - 5x + 5 = 0

OpenStudy (anonymous):

\[\LARGE Quadratic~Formula: x = \frac{-b ± \sqrt{b^2 - 4ac}}{2a}\]\[\large where,\]\[\large a = -2\]\[\large b = -5\]\[\large c = 5\]

OpenStudy (waleed_imtiaz):

\[x= -b \pm \sqrt{b^{2}-4ac}/2a\] put the values

OpenStudy (anonymous):

Lopez.. my answers dont look like that on the site

OpenStudy (anonymous):

I still dont know how to do this.:/

OpenStudy (anonymous):

\[-2x^2 - 5x + 5 = 0\]multiply by \(-1\) to make life nicer get \[2x^2 + 5x - 5 = 0\]use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=2,b=-5,c=5\]

OpenStudy (anonymous):

put the numbers in the slot,s, do the arithmetic carefully \[x=\frac{-(-5)\pm\sqrt{(-5)^2-4\times 1\times 5}}{2\times 2}\]

OpenStudy (anonymous):

click on the alternate answers the answers are there

OpenStudy (anonymous):

i got x approx -3.26556 and x approx 0.765564

OpenStudy (anonymous):

I tried to work it out and i still dont see my answer'

OpenStudy (anonymous):

 My answers look like this